# Critical Iterative Denoising A Discrete Generative Model Applied to Graphs

### v1

@@ Appendix B Generative Graph Modeling: Related Works
In parallel, several approaches address scalability challenges in graph generation (Qin et al., 2024b; Chen et al., 2023; Karami, 2024; Bergmeister et al., 2024; Boget et al., 2025). Notably, Qin et al. (2024b) and Boget et al. (2025) propose methods that enable any equivariant model to scale to large graphs. While these methods could extend our iterative denoising framework to larger graphs, we leave this exploration for future research.
Appendix C Models
C.1 GNNs architecture
@@
C.2 hyperparameters
As explained here above, the node hidden representation has size $d_{h}$ and the edge representation $d_{h}/4$. We use $d_{h}=64$ with the Qm9 dataset and $d_{h}=256$ with all the other dataets.
@@ C.3 Extra Features
Appendix D Evaluation
D.1 Molecular Benchmark
@@ D.2 Generic Graphs
Appendix E Complementary Experimental Results
@@
Appendix F Visualizations
F.1 Molecular graphs

### v2

@@ Appendix C Generative Graph Modeling: Related Works
A key limitation of current equivariant models is their reliance on dense representations and pairwise computations, which hinder scalability to large graphs. To address this, recent works have proposed scalable denoising architectures and sampling strategies (Qin et al., 2024b; Chen et al., 2023; Karami, 2024; Bergmeister et al., 2024; Boget et al., 2025). Notably, Qin et al. (2024b) and Boget et al. (2025) propose methods that enable any equivariant model to scale to large graphs. While these methods could extend our iterative denoising framework to larger graphs, we leave this exploration for future research.
Appendix D Models
D.1 GNNs architecture
@@
D.2 hyperparameters
As explained here above, the node hidden representation has size $d_{h}$ and the edge representation $d_{h}/4$. We use $d_{h}=64$ with the QM9 dataset and $d_{h}=256$ with all the other datasets.
@@ D.3 Extra Features
Appendix E Evaluation
E.1 Molecular Benchmark
@@ E.2 Generic Graphs
E.3 Computational Cost and Baselines
E.3.1 Additional Overhead from CID
CID introduces a computational overhead, as it requires training a Critic and evaluating it at each denoising step. As a first-order approximation, this doubles both training and inference time. However, the performance gain over Mask ID and Mask DDM is important, and Mask denoising models are suitable for some specific downstream applications such as molecular scaffold extension.
E.3.2 Comparison with Baselines
For both our model and baseline methods, the dominant computational cost arises from the architecture—especially from the size of edge feature vectors, on which MLPs are applied, as their number scales quadratically with the number of nodes in the graph, and linearly with the number of layers.
Below is a comparison based on DruM’s configuration files, summarizing the number of layers ($L$) and edge feature dimensions ($d_{E}$) per dataset:
Table 5: Number of layers ($L$) and hidden dimension ($d_{E}$) used for our models and DruM across datasets.
Dataset Ours ($L\times d_{E}$) DruM ($L\times d_{E}$)
Qm9 4$\times$16 8$\times$256
Zinc 4$\times$64 9$\times$128
Planar 4$\times$64 8$\times$64
SBM 4$\times$64 8$\times$64
We highlight that we evaluate our model using 500 denoising steps, whereas the baselines use 1000 steps.
In summary, our model achieves superior performance in key practical tasks, such as large molecule generation (as measured by validity and Fréchet ChemNet Distance), using fewer denoising steps, fewer layers, and smaller per-layer costs. We appreciate the reviewer’s remark to make and will update the manuscript in order to explicit these advantages. Evidence Supporting the Effectiveness of Our Method
Our work addresses a core limitation of discrete denoising models. We show that our approach consistently outperforms existing discrete diffusion methods across all evaluation metrics, including those ablating NFE’s. On this regard, we believe the paper provides sufficient evidence of our method’s effectiveness.
Appendix F Complementary Experimental Results
@@
Appendix G Visualizations
G.1 Molecular graphs